What is radius of gyration and moment of inertia of hollow sphere

Grade:12

1 Answers

Krishan Pareek
6 years ago
for hollow sphere moment of inertia = m*r*r/2for radius of gyration m*k*k = m*r*r/2k = 1/√2inertia of a hollow cylinder= inertia of a ringinertia of a solid cylinder = inertia of a disc